### v1

@@ Abstract
The burgeoning field of brain health research increasingly leverages artificial intelligence (AI) to interpret and analyze neurological data. This study introduces a novel approach towards the creation of medical foundation models by integrating a large-scale multi-modal magnetic resonance imaging (MRI) dataset derived from 41,400 participants in its own. Our method involves a novel two-stage pretraining approach using vision transformers. The first stage is dedicated to encoding anatomical structures in generally healthy brains, identifying key features such as shapes and sizes of different brain regions. The second stage concentrates on spatial information, encompassing aspects like location and the relative positioning of brain structures. We rigorously evaluate our model, BrainFounder, using the Brain Tumor Segmentation (BraTS) challenge and Anatomical Tracings of Lesions After Stroke v2.0 (ATLAS v2.0) datasets. BrainFounder demonstrates a significant performance gain, surpassing the achievements of the previous winning solutions using fully supervised learning. Our findings underscore the impact of scaling up both the complexity of the model and the volume of unlabeled training data derived from generally healthy brains, which enhances the accuracy and predictive capabilities of the model in complex neuroimaging tasks with MRI. The implications of this research provide transformative insights and practical applications in healthcare and make substantial steps towards the creation of foundation models for Medical AI. Our pretrained models and training code can be found at https://github.com/lab-smile/GatorBrain.
@@
Neuroimaging Foundation ModelSelf-Supervised LearningBrain Tumor SegmentationBrainFounder
@@
The fusion of artificial intelligence (AI) with neuroimaging, particularly magnetic resonance imaging (MRI), is forging a pivotal role in advancing brain health ( [chen_ai-based_2022],  [segato_artificial_2020],  [rao_brain_2023],  [owolabi_global_2023],  [moreno-blanco_technologies_2019],  [rajpurkar_ai_2022]  [khachaturian_accelerating_2023]). The complexity of the human brain, with its elaborate anatomy and intricate functions, poses significant challenges in neuroimaging analysis ([moor_foundation_2023], [azad_foundational_2023],[zhang_challenges_2024], [segato_artificial_2020], [rajpurkar_ai_2022]). AI’s capability to interpret complex neurological data holds the promise of enhancing diagnostic precision and deepening our understanding of brain pathology. Numerous studies have aimed to develop AI models for specific brain health analyses, each contributing to the growing body of neuroimaging research.
@@
Traditionally, neuroimaging AI models require extensive fine-tuning through supervised learning to address a specific downstream task. Modifications of the nnU-Net ([isensee_nnu-net_2021]), DeepScan ([mckinley_ensembles_2019]), and DeepMedic ([kamnitsas_efficient_2017]) architectures have performed well on a host of medical computer vision challenges such as the Brain Tumor Segmentation (BraTS) challenge ([baid_rsna-asnr-miccai_2021]), Medical Segmentation Decathlon (MSD) ([antonelli_medical_2022]), and A tumor and liver automatic segmentation challenge (ATLAS) ([Quinton2023]). Many of these advances stem from utilizing self-supervised pretraining methods on large, unlabeled datasets to transfer weights for model encoders and decoders to the smaller datasets present in the challenge ([zhou_models_2021], [tang_self-supervised_2022]). Complimentary to these pretraining modifications, there has been a recent push towards developing massive medical datasets ([mei_radimagenet_2022], [clark_cancer_2013], [bycroft_uk_2018]) to aid in the creation of these models. However, medical image analysis has yet to benefit from the recent advances in natural image analysis and language processing through models like the Segment Anything Model (SAM) ([kirillov_segment_2023]) and LLaMA ([touvron_llama_2023]). However, foundation models capable of performing few- or zero-shot learning with one set of model weights have the potential to overcome this massive fine-tuning barrier.
@@
In medical language processing, models like MI-Zero ([lu_visual_2023]) and BioViL-T ([bannur_learning_2023]) utilize contrastive learning to make significant advancements in representational analysis and zero-shot transfer learning in medical image recognition. By leveraging different learning objectives, similar image-text pairs are pulled closer in the latent space while different pairs are pushed further apart. Though such models have pushed the boundary of histopathology research and combined text-based analysis with computer vision, these models rely on having text-based prompts accompanying their training images ([tiu_expert-level_2022]).
@@
With SAM’s demonstrated success on few-shot segmentation tasks of natural images, medical models trained on purely image-based modalities have largely aimed at modifying the SAM architecture. Models like MedSAM ([ma_segment_2023]), MedLSAM ([lei_medlsam_2023]), and SAM-Med2D ([cheng_sam-med2d_2023]) focus on enhancing performance on medical segmentation tasks and bridging the gap between SAM’s generalizability on real-world images and its performance on medical tasks by adapting the SAM architecture to these medical tasks. Through a massive dataset of image-mask pairs from large medical image databases, [ma_segment_2023] crafted MedSAM for image segmentation, a technique which was further refined by including landmark localization in MedLSAM. SAM-Med2D further improved segmentation results by increasing the dataset to multiple modalities and increasing prompt density. However, these models function in 2-dimensional space, requiring 3D modalities to be sub-sampled or solved in slices ([azad_foundational_2023]). Not only is this computationally inefficient, but also the most information dense and often most valuable information is found in 3D modalities like CT or MRI. [Gong2023] aimed to address this discrepancy by adapting the SAM models to 3D space using a visual sampler and a mask decoder to aggregate layers. Their model, dubbed 3DSAM-adapter, outperformed leading segmentation models in a variety of tasks while still utilizing an algorithm that functions in 2D space, indicating that these models would benefit from the critical anatomical and spatial information found from being fully capable of functioning in 3D space.
@@
Despite the progress in medical imaging, accurately analyzing the vast a-mount of data generated by brain MRIs remains a formidable challenge ([azad_foundational_2023]). The intricate structure and function of the brain necessitate advancements in MRI analysis due to their critical impact on patient outcomes, especially in the early detection and treatment of brain disorders ([zhang_challenges_2024]). Existing AI models in neuroimaging are hampered by their need for extensive supervised learning and their limited ability to generalize across different tasks without substantial retraining, revealing a gap for a robust, adaptable model that functions in 3D space ([azad_foundational_2023], [zhang_challenges_2024]).
@@
This study presents BrainFounder, a foundational framework intended to serve as a precursor to comprehensive neuroimage foundation models. BrainFounder is designed to pave the way towards setting new standards in the accuracy and efficiency of medical AI models. We focus our study on two key tasks - brain tumor segmentation and brain lesion segmentation. A primary obstacle in creating AI models for brain tumor and brain lesion analysis is the scarcity of brain tumors within the general populace. This scarcity significantly hampers the compilation of large diseased patient datasets, which are essential for the supervised training of AI models. In response, the development process of BrainFounder incorporates a segmented approach to feature learning, specifically engineered to mitigate the challenges posed by data scarcity.
@@
In its initial phase, BrainFounder leverages an extensive dataset from brain scans of 41,400 participants. This foundational step enables the framework to effectively encode normal brain tissue structures, creating a detailed baseline of anatomical features from a predominantly healthy population. Subsequently, the framework’s training shifts focus towards identifying disease-specific attributes, such as geometric shapes of tumors and lesions and spatial placements within the brain. This dual-phase methodology significantly diminishes the extensive data requirements usually necessary for AI model training in tumor detection. Moreover, it naturally expands the dataset available for the AI to learn from in an efficient and straightforward manner, sidestepping the need for generating synthetic images. This approach mirrors the analytical techniques used by radiologists and has undergone thorough validation against the BraTS challenge and ATLAS 2.0 datasets, showcasing significant improvements over current models.
@@
BrainFounder, our novel framework, represents a pivotal advance in neuroimaging analysis by laying the groundwork for a future of comprehensive foundation models in this field. BrainFounder is designed to be adaptable for a variety of neurological tasks, including brain tumor segmentation, stroke localization, brain region segmentation, and the diagnosis of Alzheimer’s disease. By utilizing a large dataset of brain imaging from a generally healthy population, BrainFounder sets the stage for transforming clinical workflows, aiming to enhance the speed and accuracy of diagnoses across a spectrum of neurological conditions. Our methodology is twofold: firstly, leverage a broad dataset of healthy brain images to create a latent-space representation of healthy brain MRI; and secondly, introduce the ability to detect anomalies by training on anomaly specific datasets. This approach overcomes the obstacle of scarce patient data in brain tumor analysis and establishes BrainFounder as a versatile, multifaceted framework in the realm of medical diagnostics. Its extensive applicability signals a move towards more integrated and flexible approaches in neurological diagnostics.
@@
Sec. 2 Methods and Materials
2.1 Model Architecture and Pipeline
The BrainFounder framework introduces a deep learning training scheme tailored for diverse applications by showcasing a distinct approach to self-supervised pretraining followed by precise fine-tuning. This section offers a detailed examination of the framework’s architecture and its procedural pipeline and highlights the innovative phases of self-supervised pretraining, termed Stage 1 and Stage 2, before proceeding to fine-tuning for downstream tasks. The architecture and learning pathway of BrainFounder are illustrated in Figure  1. Central to BrainFounder is a vision transformer-based encoder that employs a series of self-attention mechanisms. This encoder is linked with an up-sampling decoder, tailored for segmentation tasks. This encoder based model architecture is adapted from the SwinUNETR architecture [hatamizadeh_swin_2022] with modified input channels and input hyperparemeters. BrainFounder pioneers a novel dual-phase self-supervised pretraining method, integrating self-supervised learning components within its structure. During Stage 1 pretraining, the framework is exposed to a wide-ranging dataset of brain MRIs from the UK Biobank dataset, predominantly featuring healthy individuals. This initial stage is designed to equip the model with a thorough comprehension of standard brain anatomy, utilizing self-supervised learning to enhance prediction capabilities. Stage 2 of pretraining advances the model’s proficiency by introducing it to a specialized dataset of MRIs geared towards the downstream task. This phase leverages the architecture’s refined anomaly detection skills, focusing on distinguishing deviations in brain structure.
Following the pretraining stages, BrainFounder undergoes fine-tuning on the final dataset, where the model’s encoder is enhanced through transfer learning. As depicted in Figure 1, the fine-tuning process leverages the pretrained Swin Transformer encoder from the earlier two stages. The first stage of pretraining with the UKB dataset is aimed at developing a foundational understanding of normal brain anatomy. The second stage of pretraining with unhealthy datasets builds upon this foundation by introducing pathology, thus allowing the model to learn the distinction between healthy and pathological tissues. Transfer learning is applied after each pretraining stage to retain and refine the knowledge acquired, ensuring that the model can effectively adapt to the new dataset while preserving previously learned patterns.
@@ 2.1 Model Architecture and Pipeline
In summary, the BrainFounder model’s architecture and pretraining para-digm represent a comprehensive approach to understanding and segmenting brain images, with a training pipeline that methodically builds the model’s capacity to differentiate and characterize complex patterns in MRI data. The self-supervised learning stages, coupled with the fine-tuning process, prepare BrainFounder to tackle downstream tasks with high efficiency and accuracy.
@@ 2.2 Data Acquisition and Preprocessing
Throughout our pretraining and fine-tuning, we make use of the UK Biobank (UKB), Brain Tumor Segmentation (BraTS) Challenge, and Anatomical Tracings of Lesions After Stroke v2.0 (ATLAS v2.0) datasets. The following section provides a summary of the datasets as they relate to our methodology, a summary of which can be found in Table  1.
@@
Modalities T1, T2-FLAIR T1, T1-ce, T2, T2-FLAIR T1-ce
@@ 2.2.1 UK Biobank dataset
In our first stage, we utilize T1 and T2-weighted Fluid Attenuation Inversion Recovery (T2-FLAIR) from the UK Biobank (UKB) data-set ([littlejohns_uk_2020]). These data-points were collected starting from 2014 and preprocessed by the UKB. Utilizing a comprehensive 35-minute protocol, the UKB obtained a diverse array of brain imaging modalities, including T1-weighted and T2-FLAIR structural brain MRI images ([smith_uk_2022]). We obtained all T1 and T2 images available between 2014 and 2022 from 44,172 participants with neuroimaging data. Raw T1 structural images were processed using a processing pipeline developed by UK Biobank researchers, generating additional images including segmentation between different types of matter and effectively reduced non-brain tissue interference. Volumetric measures of gray matter and internal structures were generated alongside the processed images, providing valuable insights into the characteristics of gray matter and internal structures. Each T1 structural image underwent further processing with FreeSurfer ([woolrich_bayesian_2009]), followed by a quality control check for inclusion into the data made available by UK Biobank researchers. Additionally, T2-FLAIR images were defaced to uphold participant privacy and then aligned to the corresponding T1 image, resulting in two additional product images. More volumetric measures of white matter lesions were saved as an additional data measure with both T1 and T2-FLAIR data within UK Biobank. Images were reconstructed as DICOM images and converted to the NIfTI format using dcm2niix ([li_first_2016]). Images were then defaced to preserve patient anonymity and broadcast to the MNI152 template space using FNIRT ([woolrich_bayesian_2009]). Among these 44,172 participants, 43,369 participants have both T1 and T2-FLAIR images. To build 3-dimensional foundation models of neuroimages, we selected participants with both T1 and T2-FLAIR volumes more than 100 slices, resulting in 41,400 participants, and 82,800 imaging volumes. A CONSORT diagram depicting the data used in this study can be found in Figure 3, and demographic data for participants is summarized in Figure 2. For detailed information, see the Appendix.

### v3

@@ Abstract
The burgeoning field of brain health research increasingly leverages artificial intelligence (AI) to analyze and interpret neuroimaging data. Medical foundation models have shown promise of superior performance with better sample efficiency. This work introduces a novel approach towards creating 3-dimensional (3D) medical foundation models for multimodal neuroimage segmentation through self-supervised training. Our approach involves a novel two-stage pretraining approach using vision transformers. The first stage encodes anatomical structures in generally healthy brains from the large-scale unlabeled neuroimage dataset of multimodal brain magnetic resonance imaging (MRI) images from 41,400 participants. This stage of pertaining focuses on identifying key features such as shapes and sizes of different brain structures. The second pretraining stage identifies disease-specific attributes, such as geometric shapes of tumors and lesions and spatial placements within the brain. This dual-phase methodology significantly reduces the extensive data requirements usually necessary for AI model training in neuroimage segmentation with the flexibility to adapt to various imaging modalities. We rigorously evaluate our model, BrainSegFounder, using the Brain Tumor Segmentation (BraTS) challenge and Anatomical Tracings of Lesions After Stroke v2.0 (ATLAS v2.0) datasets. BrainSegFounder demonstrates a significant performance gain, surpassing the achievements of the previous winning solutions using fully supervised learning. Our findings underscore the impact of scaling up both the model complexity and the volume of unlabeled training data derived from generally healthy brains. Both of these factors enhance the accuracy and predictive capabilities of the model in neuroimage segmentation tasks. Our pretrained models and code are at https://github.com/lab-smile/BrainSegFounder.
@@
Neuroimaging 3D Foundation Model Self-Supervised Learning Brain Tumor Segmentation Multi-modal MRI
@@
The fusion of artificial intelligence (AI) with neuroimaging analysis, particularly multimodal MRI, is forging a pivotal role in advancing brain health ([chen_ai-based_2022], [segato_artificial_2020], [rao_brain_2023], [owolabi_global_2023], [moreno-blanco_technologies_2019], [rajpurkar_ai_2022] [khachaturian_accelerating_2023]). The complexity of the human brain, with its elaborate anatomy and intricate functions, poses significant challenges in neuroimaging analysis ([moor_foundation_2023], [azad_foundational_2023],[zhang_challenges_2024], [segato_artificial_2020], [rajpurkar_ai_2022]). AI’s capability to interpret complex neurological data has the potential to enhance diagnostic precision and deepen our understanding of brain pathology. Numerous studies have aimed to develop AI models for specific brain health analyses, each contributing to the growing body of neuroimaging research.
@@
Traditionally, neuroimaging AI models require extensive fine-tuning through supervised learning to address a specific downstream task. Modifications of the nnU-Net ([isensee_nnu-net_2021]), DeepScan ([mckinley_ensembles_2019]), and DeepMedic ([kamnitsas_efficient_2017]) architectures have performed well on a host of medical computer vision challenges such as the Brain Tumor Segmentation (BraTS) challenge ([baid_rsna-asnr-miccai_2021]), Medical Segmentation Decathlon (MSD) ([antonelli_medical_2022]), and A tumor and liver automatic segmentation challenge (ATLAS) ([Quinton2023]). Many of these advances stem from utilizing self-supervised pretraining methods on large, unlabeled datasets to transfer weights for model encoders and decoders to the smaller datasets present in the challenge ([zhou_models_2021], [tang_self-supervised_2022]). Complementary to these pretraining modifications, there has been a recent push towards developing massive medical datasets ([mei_radimagenet_2022], [clark_cancer_2013], [bycroft_uk_2018]) to aid in the creation of these models. However, medical image analysis has yet to benefit from the recent advances in natural image analysis and language processing through models like the Segment Anything Model (SAM) ([kirillov_segment_2023]) and LLaMA ([touvron_llama_2023]).
@@
In medical language processing, models like MI-Zero ([lu_visual_2023]) and BioViL-T ([bannur_learning_2023]) utilize contrastive learning to make significant advancements in representational analysis and zero-shot transfer learning in medical image recognition. By leveraging different learning objectives, similar image-text pairs are pulled closer in the latent space while dissimilar pairs are pushed further apart. Such models have pushed the boundary of histopathology research and combined text-based analysis with computer vision. Yet, they rely on having text-based prompts accompanying their training images ([tiu_expert-level_2022]).
@@
With SAM’s demonstrated success on few-shot segmentation tasks of natural images, recent research into medical image segmentation models has primarily modified the SAM architecture. Models like MedSAM ([ma_segment_2023]), MedLSAM ([lei_medlsam_2023]), and SAM-Med2D ([cheng_sam-med2d_2023]) focus on bridging the gap between SAM’s generalizability on real-world images and its performance on medical tasks. They accomplish this by adapting the SAM architecture to these medical tasks. [ma_segment_2023] crafted MedSAM for image segmentation by constructing a massive dataset of image-mask pairs derived from sizeable medical image databases. MedLSAM further refined upon MedSAM by including landmark localization. SAM-Med2D further improved segmentation results by increasing the dataset to multiple modalities and increasing prompt density. However, these models function in 2-dimensional space, requiring 3-dimensional (3D) modalities to be sub-sampled or solved in slices ([azad_foundational_2023]). Not only is this computationally inefficient, but the most dense and often most valuable information is found in 3D modalities like CT or MRI. [Gong2023] aimed to address this discrepancy by adapting the SAM models to 3D space using a visual sampler and a mask decoder to aggregate layers. Their model, dubbed 3DSAM-adapter, outperformed leading segmentation models in various tasks while still utilizing an algorithm that functions in 2D space. These results indicated that these models would benefit from the critical anatomical and spatial information found from being fully capable of functioning in 3D space.
@@
Despite the progress in medical imaging, holistically analyzing the vast amount of data generated by brain MRIs remains a formidable challenge ([azad_foundational_2023]). The intricate structure and function of the brain necessitate advancements in MRI analysis due to their critical impact on patient outcomes, especially in the early detection and treatment of brain disorders ([zhang_challenges_2024]). Existing AI models in neuroimaging are hampered by their need for extensive supervised learning and their limited ability to generalize across different tasks without substantial retraining, revealing a gap for a robust, adaptable model that functions in 3D space ([azad_foundational_2023], [zhang_challenges_2024]).
@@
This study presents BrainSegFounder, a 3D foundational framework for multimodal neuroimage segmentation. BrainSegFounder is designed to pave the way towards setting new standards for the accuracy and efficiency of medical AI models. We focus our study on two essential tasks - brain tumor segmentation and brain lesion segmentation. A primary obstacle in creating AI models for brain tumor and brain lesion analysis is the scarcity of brain tumors within the general population. This scarcity significantly hampers the compilation of large-scale diseased patient datasets, which are essential for the supervised training of AI models. In response, the development process of BrainSegFounder incorporates a multi-stage approach to feature learning, specifically engineered to mitigate the challenges posed by data scarcity.
@@
In its initial phase, BrainSegFounder leverages an extensive dataset from brain scans of 41,400 participants from the United Kingdom. This foundational step enables the framework to effectively encode generally healthy brain tissue structures, creating a detailed baseline of anatomical features from a predominantly healthy population. Subsequently, the framework’s training shifts focus towards identifying disease-specific attributes, such as geometric shapes of tumors and lesions and spatial placements within the brain. This dual-phase methodology significantly diminishes the extensive data requirements usually necessary for AI model training in tumor detection. Moreover, it naturally expands the dataset available for the AI to learn from efficiently and straightforwardly, sidestepping the need for generating synthetic images. This approach mirrors the analytical techniques used by radiologists and has undergone thorough validation against the BraTS challenge and ATLAS 2.0 datasets, showcasing significant improvements over current models.
@@
BrainSegFounder, our novel framework, represents a pivotal advance in neuroimaging analysis by laying the groundwork for a future of comprehensive foundation models in this field. BrainSegFounder is designed to be adaptable for various neurological tasks, including brain tumor segmentation, stroke localization, brain region segmentation, and the diagnosis of Alzheimer’s disease. By utilizing a large dataset of brain imaging from a generally healthy population, BrainSegFounder sets the stage for transforming clinical workflows, aiming to enhance the speed and accuracy of diagnoses across a spectrum of neurological conditions. The contribution of this work is twofold: 1) BrainSegFounder leverages a large-scale multi-modal 3D neuroimaging dataset of generally healthy brain images to create a latent-space representation of healthy brain MRI; 2) it gains the ability to detect anomalies by training on anomaly-specific datasets with adaptability between imaging modalities. This approach effectively addresses the challenge of limited patient data in diseased neuroimaging analysis, establishing BrainSegFounder as a versatile framework in medical diagnostics. Its broad applicability signifies a shift towards more integrated and adaptable methodologies in neurological diagnostics.
@@
The BrainSegFounder framework introduces a deep learning training scheme tailored for diverse applications by showcasing a distinct approach to self-supervised pretraining followed by precise fine-tuning. This section offers a detailed examination of the framework’s architecture and its procedural pipeline. It highlights the multi-stage self-supervised pretraining, termed Stage 1 and Stage 2, before proceeding to fine-tuning for downstream tasks. Figure  1 illustrates BrainSegFounder’s architecture. Central to BrainSegFounder is a vision transformer-based encoder that employs a series of self-attention mechanisms. This encoder is linked with an up-sampling decoder tailored for segmentation tasks. The architecture is adapted from the SwinUNETR architecture [hatamizadeh_swin_2022] with modified input channels and input hyperparameters. BrainSegFounder pioneers a novel dual-phase self-supervised pretraining method, integrating self-supervised learning components within its structure. Stage 1 pretraining exposes the framework to a wide-ranging dataset of brain MRIs from the UK Biobank dataset, predominantly consisting of healthy individuals. This initial stage equips the model with a thorough comprehension of standard brain anatomy, utilizing self-supervised learning to enhance prediction capabilities. Stage 2 of pretraining advances the model’s proficiency by introducing it to a specialized MRI dataset geared toward the downstream task. This phase leverages the architecture’s refined anomaly detection skills, focusing on distinguishing deviations in brain structure.
Following pretraining, BrainSegFounder undergoes fine-tuning on the final dataset, where transfer learning enhances the model’s encoder. As depicted in Figure 1, the fine-tuning process leverages the pretrained Swin Transformer encoder from the earlier two stages. The first pretraining stage on the UKB dataset develops a foundational understanding of normal brain anatomy. The second stage of pretraining with diseased datasets builds upon this foundation by introducing pathology, thus allowing the model to learn the distinction between healthy and pathological tissues. Transfer learning is applied after each pretraining stage to retain and refine the knowledge acquired, ensuring that the model can effectively adapt to the new dataset while preserving previously learned patterns.
@@
In summary, the BrainSegFounder model’s architecture and pretraining paradigm represent a comprehensive approach to understanding and segmenting brain images, with a training pipeline that methodically builds the model’s capacity to differentiate and characterize complex patterns in 3D MRI data without external annotation. Together, the self-supervised learning stages and the fine-tuning process prepare BrainSegFounder to tackle downstream tasks with high efficiency and accuracy.
@@ 1.1 Data Acquisition and Preprocessing
Throughout our pretraining and fine-tuning, we make use of the UK Biobank (UKB), Brain Tumor Segmentation (BraTS) Challenge, and Anatomical Tracings of Lesions After Stroke v2.0 (ATLAS v2.0) datasets. The following section summarizes the dataset information that is pertinent to our study. Table  1 provides an overview of this information.
@@
Modalities T1w, T2-FLAIR T1w, T1-ce, T2w, T2-FLAIR T1-ce
@@ 1.1.1 UK Biobank dataset
In our first stage, we utilize T1-weighted (T1w) and T2-weighted Fluid Attenuation Inversion Recovery (T2-FLAIR) from the UK Biobank (UKB) dataset ([littlejohns_uk_2020]). These data points were collected starting from 2014 and preprocessed by the UKB. Utilizing a comprehensive 35-minute protocol, the UKB obtained many brain imaging modalities, including T1w and T2-FLAIR structural brain MRI images ([smith_uk_2022]). We obtained all T1w and T2-FLAIR images available between 2014 and 2022 from 44,172 participants with neuroimaging data. Raw T1w structural volumes were processed using a processing pipeline developed by UK Biobank researchers that consisted primarily of tools from FSL and Freesurfer. The pipeline generated additional images like segmentations between different types of matter and effectively reduced the non-brain tissue interference. Volumetric measures of gray matter and internal structures were generated alongside the processed images, providing valuable insights into the characteristics of gray matter and internal structures. Each T1w structural image underwent further processing with FreeSurfer ([woolrich_bayesian_2009]), followed by a quality control check for inclusion into the data made available by UK Biobank researchers. Additionally, T2-FLAIR images were aligned to the corresponding T1 image, resulting in two additional product images. The UK Biobank saves volumetric measures of white matter lesions as additional data with both T1w and T2-FLAIR volumes. Images were reconstructed as DICOM images and converted to the NIfTI format using dcm2niix ([li_first_2016]). All imaging volumes were defaced to preserve participant anonymity and broadcast to the MNI152 template space using FNIRT ([woolrich_bayesian_2009]). Among these 44,172 participants, 43,369 participants have both T1w and T2-FLAIR images. To build 3D foundation models of neuroimages, we selected participants who had at least 100 slices in both T1w and T2-FLAIR volumes. This criteria resulted in 41,400 participants and 82,800 imaging volumes. A CONSORT diagram depicting the data used in this study can be found in Figure 3, and demographic data for participants is summarized in Figure 2. For detailed information, see the Appendix.